# Singularity of the axisymmetric stagnation-point-like solution within a cylinder of the 3D Euler incompressible fluid equations

### v1

@@ 4 Conclusion
We have applied the vorticity-stretching ansatz by Gibbon et al. (1999) on cylindrical domains with boundaries, to explore the potential formation of singularities in the three-dimensional incompressible Euler equations. The study incorporates axisymmetry, a cylindrical spatial domain, and no-flow boundary conditions. By employing the Lagrangian framework, explicit solutions for the planar vorticity, its stretching rate, and fluid pathlines are obtained through the use of conserved quantities—constants of motion—that remain invariant along particle trajectories. A key new result stemming from our axisymmetry assumption is that explicit solutions for the Eulerian velocity components can be obtained.
@@
The analysis reveals that the key factor determining whether or not a singularity forms in finite time is the local behavior of the initial vorticity stretching rate near the position of its minimum. Specifically, the support of this minimum (whether it is at the centre point or at a ring) and the way the vorticity stretching rate behaves in its vicinity dictate the development of blowup in the solution. These are considered local properties of the initial data and do not depend on the global configuration of the flow.
@@
Further investigation into the asymptotic behavior of the system shows that different power-law approximations of the initial vorticity stretching rate near the position of its minimum lead to distinct types of singularity. We identified two critical thresholds: one marking the transition from regular to singular solutions (in terms of the formula for the singularity time $T_{*}$), and another distinguishing different rates at which quantities converge or diverge near the singularity time.
@@
The emergence of a ring-shaped blowup region at the “corner” boundary of a cylinder brings to mind the numerical observations in Luo & Hou (2014), where a ring-localised singularity was detected in the 3D axisymmetric Euler equations. Our theoretical framework, although in a different class of solutions with respect to the reflection $z\to-z$ (in our case both $u_{r}$ and $u_{\theta}$ are even functions of $z$, while in Luo and Hou only $u_{r}$ is even in $z$, while $u_{\theta}$ is odd in $z$), allows for ring-shaped minimum profiles of $\gamma_{0}(r_{0})$, provides analytical support for such non-central singularities and explains the lower blowup threshold in this configuration due to geometric inhibition. Interestingly, there is a work in the literature that belongs to the class studied in our paper, and therefore can be discussed quantitatively in terms of our results: the paper by Ohkitani and Gibbon (Ohkitani & Gibbon, 2000, Section II), where two ring-like solutions were computed numerically: Initial condition 2 (in Section II B) and Initial condition 3 (in Section II C). Initial condition 2 corresponds to a non-central minimum with $n=2$, while Initial condition 3 corresponds to a central minimum with $n=2$. According to our theoretical framework, for Initial condition 2 the exponent $n$ lies at the blowup threshold (hence regular behavior is expected), while for Initial condition 3 the exponent $n$ lies in the blowup-rate threshold (hence singular behaviour is expected). This prediction is in agreement with the numerical results reported in their study: compare figure 8 (for Initial condition 2), showing a tame behaviour, with figure 11 (for Initial condition 3), showing a singular behaviour. This consistency between our theoretical predictions and prior numerical simulations underscores the validity and predictive power of the proposed framework in characterizing singularity formation in such systems.
@@
Our work not only provides a platform for studying the fundamental question of regularity in three-dimensional Euler flows but also serves as a mathematical illustration of vortex interactions under Gibbon et al.’s model. It highlights how local geometric features of the initial conditions can drive the formation of singularities, offering insight into the mechanisms underlying turbulent behavior in idealised flows.
@@ 5 Acknowledgements
This publication has emanated from research supported in part by a grant from Taighde Éireann – Research Ireland under Grant number 18/CRT/6049. For the purpose of Open Access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.

### v2

@@ 4 Conclusion
We have applied the vorticity-stretching ansatz by Gibbon et al. (1999) on cylindrical domains with boundaries, to explore the potential formation of singularities in the three-dimensional incompressible Euler equations. The study incorporates axisymmetry, a cylindrical spatial domain, and no-flow boundary conditions. By employing the Lagrangian framework, explicit solutions for the planar vorticity, its stretching rate, and fluid pathlines are obtained through the use of conserved quantities—constants of motion—that remain invariant along particle trajectories. A key new result stemming from our axisymmetry assumption is that explicit solutions for the Eulerian velocity components can be obtained under specific initial stretching rate, such as the parabolic initial $\gamma_{0}(r_{0})$ discussed in Section 3.1.
@@
The analysis reveals that the key factor determining whether or not a singularity forms in finite time is the local behavior of the initial vorticity stretching rate near the position of its minimum. Specifically, the position of this minimum (whether it is at the center point or at a ring) and the way the initial vorticity stretching rate behaves in its vicinity dictate the development of blowup in the solution. These are considered local properties of the initial data and do not depend on the global configuration of the flow. An exception to this result occurs when the minimum of $\gamma$ is initially attained at several places. Then, the structure of the minimum at each of these places needs to be considered in order to determine which one dominates, based on the asymptotic behavior of the corresponding auxiliary integrals $\Lambda_{1}$ defined in equation (64).
@@
Further investigation into the asymptotic behavior of the system shows that different power-law approximations of the initial vorticity stretching rate near the position of its minimum lead to distinct types of singularity. We identified two critical thresholds: one marking the transition from regular to singular solutions (in terms of the formula for the singularity time $T_{*}$), and another one distinguishing different rates of blowup or vanishing near the singularity time.
@@
The emergence of a ring-shaped blowup region at the “corner” boundary of a cylinder brings to mind the numerical observations in Luo and Hou (2014), where a ring-localized singularity was detected in the 3D axisymmetric Euler equations. Our theoretical framework, although in a different class of solutions with respect to the reflection $z\to-z$ (in our case both $u_{r}$ and $u_{\theta}$ are even functions of $z$, while in Luo and Hou only $u_{r}$ is even in $z$, while $u_{\theta}$ is odd in $z$), allows for ring-shaped minimum profiles of $\gamma_{0}(r_{0})$, provides analytical support for such non-central singularities and explains the lower blowup threshold in this configuration due to geometric inhibition.
Interestingly, there is a work in the literature that belongs to the class studied in our paper, and therefore can be discussed quantitatively in terms of our results: the paper by Ohkitani and Gibbon (Ohkitani and Gibbon, 2000, Section II), where two ring-like solutions were computed numerically: Initial condition 2 (in Section II B) and Initial condition 3 (in Section II C). Initial condition 2 corresponds to a non-central minimum with $n=2$, while Initial condition 3 corresponds to a central minimum with $n=2$. According to our theoretical framework, for Initial condition 2 the exponent $n$ lies at the blowup threshold (hence regular behavior is expected), while for Initial condition 3 the exponent $n$ lies in the blowup-rate threshold (hence singular behavior is expected). This prediction is in agreement with the numerical results reported in their study: compare figure 8 (for Initial condition 2), showing a tame behavior, with figure 11 (for Initial condition 3), showing a singular behavior. This consistency between our theoretical predictions and prior numerical simulations underscores the validity and predictive power of the proposed framework in characterizing singularity formation in such systems.
@@
Our work not only provides a platform for studying the fundamental question of regularity in three-dimensional Euler flows but also serves as a mathematical illustration of vortex interactions under Gibbon et al.’s model. It highlights how local geometric features of the initial conditions can drive the formation of singularities, offering insight into the mechanisms underlying turbulent behavior in idealized flows.
@@ 5 Acknowledgements
This publication has emanated from research supported in part by a grant from Taighde Éireann – Research Ireland under Grant number 18/CRT/6049. For the purpose of Open Access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission. AI tools were used for translation and language editing; all scientific content, analysis, and conclusions were produced by the authors.